OpenStudy (anonymous):

solve by elimination 2r-5s=1 and 5r+2s=46

OpenStudy (anonymous):

2r -5s =1 ----------(i) 5r +2s = 46 ---------(ii) (i)*2 is 4r -10s = 2 (ii)*5 is 25r +10s = 230 Add 29r = 232 Divide by 29 r = 8 2r -5s =1 2*8 -5s =1 16 -5s = 1 16 -1 = 5s 15 = 5s 3 = s Answer r = 8 and s = 3

OpenStudy (anonymous):

r= (1+5s)/2 46= 5[(1+5s)/2] +2s s=3 r=8

OpenStudy (anonymous):

-5(2r-5s)=1 distribute so -10r+25s=-5 2(5r+2s)=46 distribute so 10r+4s=92 29s=87, s=3. plug in to equation to get value of r